OpenStudy (anonymous):

Weighing the baby was a problem. So i held the baby while the nurse read of 78 kg on the scale. Then the nurse held the baby while i read off 69 kg. I then held the nurse, while the baby read off 137 kg. What is the combined weight of all three people?

OpenStudy (saifoo.khan):

You have a baby? :O

OpenStudy (anonymous):

137+9 right? 146

OpenStudy (anonymous):

nope. math, problem

OpenStudy (anonymous):

i was wrong? QQ

OpenStudy (anonymous):

no i meant i don't have a baby (;

OpenStudy (saifoo.khan):

LOL.

OpenStudy (anonymous):

oh, lol

OpenStudy (anonymous):

Baby + Me = 78 Nurse + Baby = 69 Me + Nurse = 137 Nurse = 137 - Me 137 - me + baby = 69 Me = 137 - 69 + baby Me = 68 + Baby Baby + Baby + 68 = 78 Baby = 5

OpenStudy (anonymous):

Me = 73 Nurse = 64 Baby = 5 Total is 142

OpenStudy (anonymous):

THANK YOU <3

OpenStudy (anonymous):

"Good Answer" and +Fan's welcomed :P